A 0.020M solution of the weak acid HA is 0.60% ionized. Calculate the value of the ionization constant for this acid.

1
Expert's answer
2021-12-03T02:42:01-0500

"100-0.60=99.4" %


"\frac{99.4×0.020}{100}=0.0199"


Constant "=\frac{0.0199^2}{0.020-0.0199}=3.96"